OpenStudy (barreraa):

Simplify: –4(x – 6) + 3(a – 7) (Points : 1) –4x + 3a – 45 –4x + 3a + 45 –4x + 3a – 3 –4x + 3a + 3

OpenStudy (anonymous):

–4(x – 6) + 3(a – 7) We need to distribute to simplify. -4(x - 6) Starting with the first part. -4 * x = -4x -4 * -6 = 24 ( 2 negatives = a positive) So, the first part is -4x + 24. 3(a - 7) Second part now. 3 * a = 3a 3 * -7 = -21 Second part is 3a - 21. Putting them together. -4x + 24 + 3a - 21 Combine like terms. -4x + 3a + 3 <--- Answer.